# Deep Stochastic Kinematic Models for Probabilistic Motion Forecasting in Traffic

### v1

@@ Abstract
Kinematic priors have shown to be helpful in boosting generalization and performance in prior work on trajectory forecasting. Specifically, kinematic priors have been applied such that models predict a set of actions instead of future output trajectories. By unrolling predicted trajectories via time integration and models of kinematic dynamics, predicted trajectories are not only kinematically feasible on average but also relate uncertainty from one timestep to the next. With benchmarks supporting prediction of multiple trajectory predictions, deterministic kinematic priors are less and less applicable to current models. We propose a method for integrating probabilistic kinematic priors into modern probabilistic trajectory forecasting architectures. The primary difference between our work and previous techniques is the analytical quantification of variance, or uncertainty, in predicted trajectories. With negligible additional computational overhead, our method can be generalized and easily implemented with any modern probabilistic method that models candidate trajectories as Gaussian distributions. In particular, our method works especially well in unoptimal settings, such as with small datasets or in the presence of noise. Our method achieves up to a 50% performance boost in small dataset settings and up to an 8% performance boost in large-scale learning compared to previous kinematic prediction methods on SOTA trajectory forecasting architectures out-of-the-box, with minimal fine-tuning. In this paper, we show four analytical formulations of probabilistic kinematic priors which can be used for any Gaussian Mixture Model (GMM)-based deep learning models, quantify the error bound on linear approximations applied during trajectory unrolling, and show results to evaluate each formulation in trajectory forecasting.
@@ I Introduction
Rapid advancements in deep learning research directly support improvements on learning-based tasks in autonomous driving. In the last five years, perception and planning models for autonomous driving have not only seen great progress but also scaled larger in size, bringing state-of-the-art (SOTA) farther from portability and resource efficiency. For example, NVIDIA’s end-to-end driving network [1] from 2016 contained about 250k parameters on a convolutional neural network (CNN) architecture, while vision transformer (ViT)-based SOTA end-to-end driving model TransFuser [2], published in 2022, has over 168 million parameters! Similar patterns have emerged for other autonomous driving tasks, such as trajectory forecasting and small-scale agent simulation, where model size and complexity continue to grow. As model sizes become bigger and performance increases, so does the complexity of learning the basics of vehicle dynamics, where, most of the time, vehicles travel in relatively straight lines behind the vehicle directly in front.
@@
Figure 1: Motivating example for probabilistic kinematic priors. In real world traffic, vehicles have a constrained range of behaviors. For instance, it is not possible for a vehicle to move side to side directly, and driving in reverse is highly unlikely. Without kinematic priors, neural networks may search the space of all trajectories, possible or impossible. Without accounting for analytical variances in trajectories as in previous work, the range of possible future trajectories may also be unrealistic. A vehicle’s immediate trajectory milliseconds into the future may have much less uncertainty than few seconds into the future.
@@
Since training a deep neural network is costly, it is beneficial on both resource consumption and technique generalization to incorporate the use of existing dynamics models in the training process, combining existing knowledge with powerful modern architectures. Kinematic models have already been widely used in various autonomous driving tasks, especially with regards to trajectory forecasting and simulation tasks [3, 4, 5, 6]. These models explicitly describe how changes in the input parameters influence the output of the dynamical system. Typically, kinematic input parameters are often provided by either the robot policy as an action, or by a human in direct interaction with the robot, e.g. steering, throttle, and brake for driving a vehicle. For tasks modeling decision-making, such as trajectory forecasting of traffic agents, modeling the input parameters may be more descriptive and interpretable than modeling the output directly. Moreover, kinematic models relate the input actions directly to the output observation; thus, any output of the kinematic model should, at the very least, be physically feasible in the real world [3].
@@
Figure 2: Qualitative example. We visualize an example of the mean trajectory of the highest scored Gaussian from the results of Table II. Standard deviations are visualized as ellipses at each second into the future, and ground truth trajectories are drawn in green. Our method (blue) not only predicts a smoother, less jagged mean trajectory, but also provides more realistic spread of trajectories into the future. The baseline method (red) shows uneven speeds (ellipses are at uneven intervals) in addition to uncertainty extending far beyond the road boundaries.
@@
While previous works already use kinematic models to provide feasible output trajectories, none of them consider compounding uncertainty into the time horizon analytically. Current works apply kinematic models deterministically, either for deterministic trajectory prediction, or solely to predict the average trajectory in probabilistic settings. However, uncertainty (or variances) can also propagate naturally throughout timesteps via kinematic relationships, which are not explicitly modeled in previous works. For probabilistic methods, the uncertainty is typically handled as a learnable matrix, and the kinematic model is used for unrolling the average trajectory predictions.
@@
Kalmin filtering is another relevant, classical technique that uses kinematic models for trajectory prediction. Kalman filters excel in time-dependent settings where signals may be contaminated by noise, such as object tracking and motion prediction [7, 8]. Unlike classical Kalman filtering methods for motion prediction, however, we focus on models of learnable actions and behaviors, rather than fixed, constant models of acceleration and velocity. We believe that, with elements from classical methods such as Kalman filtering in addition to powerful modern deep learning architectures, motion prediction in traffic can have the best of both worlds where modeling capability meets stability.
In this paper, we go one step further from previous work by considering the relationship between distributions of kinematic parameters to distributions of trajectory rollouts, rather than single deterministic kinematic parameter predictions to the mean of trajectory rollouts [3, 6]. We hypothesize that modeling kinematic parameters probabilistically will capture how human decision making uncertainty reflects in the trajectory uncertainty more naturally, resulting in both better performance, realistic trajectories, and faster learning. In our results, we observe unnatural progression of uncertainty through time when standard deviation across time steps are treated independently and as learnable parameters. We aim to resolve this by defining the analytical relationship between uncertainty in the kinematic input parameters and the uncertainty in Euclidean trajectory rollouts.
We present a method for incorporating car-following models as priors into modern trajectory forecasting neural network architectures. By leveraging the advantages of both network complexity and simplicity of kinematic relationships, the network can prioritize learning the complex patterns of human behavior on top of basic vehicle dynamics. Additionally, we present results on four different kinematic formulations, which provide insights on differences between different kinematic representations for learning.
@@ II-A Trajectory Forecasting for Traffic
Traffic trajectory forecasting is a popular task where the goal is to predict the short-term future trajectory of multiple agents in a traffic scene. Being able to predict the future positions and intents of each vehicle provide context for other modules in autonomous driving, such as path planning. Large, robust benchmarks such as the Waymo Motion Dataset [9, 10], Argoverse [11], and the NuScenes Dataset [12] have provided a standardized setting for advancements in the task, with leaderboards showing clear rankings for state-of-the-art models. Amongst the top performing architectures, most are based on Transformers for feature extraction [13, 14, 15, 16, 17]. Current SOTA models also model trajectory prediction probabilistically, as inspired by the use of GMMs in Multipath [18].
@@ V-B Performance on a Smaller Dataset Setting
We train the baseline model and all formulations on only 1% of the original Waymo dataset and benchmark their performance on 100% of the evaluation set in Table II. All experiments were trained over 50 epochs. In the small dataset setting, we observe that providing a kinematic prior in any form improves performance for minimum final displacement error (minFDE). Additionally, we observe better performance across all metrics for formulations 1, 2, and 1 with interpolation. Figure 4 shows how all kinematic formulations improve convergence speed over the baseline, with Formulation 4 (acceleration and steering) converging most quickly. Overall, Formulation 1 provides the greatest boost in mAP performance, with over $13\%$ gain over the vanilla baseline and $50\%$ over the deterministic baseline DKM. Additionally, Formulation 4 provides a $12\%$ reduction in minADE, while Formulation 3 provides a $50\%$ reduction in minFDE and a $6\%$ reduction in Miss Rate. In general, all formulations provide similar benefits in the small dataset setting, with the most general performance benefit coming from Formulation 3. While Formulation 4 also provides a comparable performance boost, the difference from Formulation 3 may be attributed to compounding error from second-order approximation.
Compared to the results from Table I, the effects of kinematic priors in learning are much more pronounced. Since kinematic priors naturally relate the position at one timestep to the position at the next, we believe that the performance boost can be attributed to this natural relationship. In backpropagation, optimization of one position further into the time horizon directly influences predicted positions at earlier timesteps via the kinematic model. Without the kinematic prior, the relation between timesteps may be implicitly related through neural network parameters. When the model lacks data to form a good model of how an agent moves through space, the kinematic model can compensate to model such simple relationships.
@@ V-C Performance in the Presence of Noise
TABLE III: Percentage (%) degradation of performance in the presence of noise. We also compare the robustness of each method by measuring the impact on performance in the presence of noise. In the table, we compute the % change in performance between perturbed evaluation and clean evaluation, relative to each method. We observe that the model trained with Formulation 4 (steering and acceleration representations) offers the greatest robustness over other formulations in the presence of noisy inputs.
@@ VI Discussion and Conclusion
We also observe that when there is large-scale data to learn a good model of how vehicles move, the effects of kinematic priors are less pronounced. This can be observed from the less obvious improvements over the baseline in Table I compared to Table II. We conjecture that model complexity and dataset size will eventually out-scale the effects of the kinematic prior. With enough resources and high-quality data, trajectory forecasting models will learn to “reinvent the steering wheel”, or implicitly learn how vehicles move via the complexity of the neural network. For practical autonomous systems, portability and adaptability are important problems to consider aside into the future. We hope that kinematic priors can be considered in trajectory forecasting and simulation as deployment ventures past well-represented urban scenarios to suboptimal settings, whether it is with noisy sensors, smaller datasets to adapt from, or downscaled models for inference speed and portability.

### v2

@@ Abstract
In trajectory forecasting tasks for traffic, future output trajectories can be computed by advancing the ego vehicle’s state with predicted actions according to a kinematics model. By unrolling predicted trajectories via time integration and models of kinematic dynamics, predicted trajectories should not only be kinematically feasible but also relate uncertainty from one timestep to the next. While current works in probabilistic prediction do incorporate kinematic priors for mean trajectory prediction, variance is often left as a learnable parameter, despite uncertainty in one time step being inextricably tied to uncertainty in the previous time step. In this paper, we show simple and differentiable analytical approximations describing the relationship between variance at one timestep and that at the next with the kinematic bicycle model. These approximations can be easily incorporated with negligible additional overhead into any existing trajectory forecasting framework utilizing probabilistic predictions, whether it is autoregressive or one-shot prediction. In our results, we find that encoding the relationship between variance across timesteps works especially well in unoptimal settings, such as with small or noisy datasets. We observe up to a 50% performance boost in partial dataset settings and up to an 8% performance boost in large-scale learning compared to previous kinematic prediction methods on SOTA trajectory forecasting architectures out-of-the-box, with no fine-tuning. In this paper, we show four analytical formulations of probabilistic kinematic priors which can be used for any Gaussian Mixture Model (GMM)-based deep learning models, quantify the error bound on linear approximations applied during trajectory unrolling, and show results to evaluate each formulation in trajectory forecasting.
@@ I Introduction
Motion forecasting in traffic involves predicting the next several seconds of movement for select actors in a scene, given the context of historical trajectories and the local environment. There are several different approaches to motion forecasting in deep learning, but one thing is in common—scaling resources, dataset sizes, and model sizes will always help generalization capabilities. However, the difficult part of motion forecasting is not necessarily how the vehicles move, but rather, figuring out the why, as there is always a human (at least, for now) behind the steering wheel.
@@
Figure 1: Motivating example for probabilistic kinematic priors. In real-world traffic, vehicles have a constrained range of behaviors. For instance, it is not possible for a vehicle to move side to side directly, and driving in reverse is highly unlikely. Without kinematic priors, neural networks may search the space of all trajectories, possible or impossible. Without accounting for analytical variances in trajectories as in previous work, the range of possible future trajectories may also be unrealistic. A vehicle’s immediate trajectory milliseconds into the future may have much less uncertainty than a few seconds into the future.
As model sizes become bigger and performance increases, so does the complexity of learning the basics of vehicle dynamics, where, most of the time, vehicles travel in relatively straight lines behind the vehicle leading directly in front. Since training a deep neural network is costly, it is beneficial for both resource consumption and model generalization to incorporate the use of existing dynamics models in the training process, combining existing knowledge with powerful modern architectures. With embedded priors describing the relationships between variables across time, perhaps a network’s modeling power can be better allocated to understanding the why’s of human behavior, rather than the how’s.
@@
Kinematic models have already been widely used in various autonomous driving tasks, especially trajectory forecasting and simulation tasks [1, 2, 3, 4]. These models explicitly describe how changes in the input parameters influence the output of the dynamical system. Typically, kinematic input parameters are often provided by either the robot policy as an action, or by a human in direct interaction with the robot, e.g. steering, throttle, and brake for driving a vehicle. For tasks modeling decision-making, such as trajectory forecasting of traffic agents, modeling the input parameters may be more descriptive and interpretable than modeling the output directly. Moreover, kinematic models relate the input actions directly to the output observation; thus, any output of the kinematic model should, at the very least, be physically feasible in the real world [1].
@@
Figure 2: Qualitative example. We visualize an example of the mean trajectory of the highest-scored Gaussian from the results of Table II. Standard deviations are visualized as ellipses at each second into the future, and ground truth trajectories are drawn in green. Our method (blue) not only predicts a smoother, less jagged mean trajectory but also provides a more realistic spread of trajectories into the future. The baseline method (red) shows uneven speeds (ellipses are at uneven intervals) in addition to uncertainty extending far beyond the road boundaries.
@@
While previous works already use kinematic models to provide feasible output trajectories, none consider compounding uncertainty into the time horizon analytically. Previous works apply kinematic models deterministically, either for deterministic trajectory prediction or to predict the average trajectory in probabilistic settings, leaving the variance as a learnable matrix.
@@
Kalman filtering is a relevant and classical technique that uses kinematic models for trajectory prediction. Kalman filters excel in time-dependent settings where signals may be contaminated by noise, such as object tracking and motion prediction [5, 6]. Unlike classical Kalman filtering methods for motion prediction, however, we focus on models of learnable actions and behaviors, rather than constant models of fixed acceleration and velocity. We believe that, with elements from classical methods such as Kalman filtering in addition to powerful modern deep learning architectures, motion prediction in traffic can have the best of both worlds where modeling capability meets stability.
In this paper, we go one step further from previous work by considering the relationship between distributions of kinematic parameters to distributions of trajectory rollouts, rather than single deterministic kinematic parameter predictions to the mean of trajectory rollouts [1, 4]. We hypothesize that modeling uncertainty explicitly across timesteps according to a kinematic model will result in better performance, realistic trajectories, and stable learning, especially in disadvantageous settings such as small or noisy datasets, which can be common settings, especially in fine-tuning. We run experiments on four different kinematic formulations with Motion Transformer (MTR) [7] and observe up to a 50% performance boost in partial-dataset settings and up to an 8% performance boost in full dataset settings on mean average precision (mAP) metrics for the Waymo Open Motion Dataset (WOMD) [8] compared to with kinematic priors enforced on mean trajectories only.
@@ II-A Trajectory Forecasting for Traffic
Traffic trajectory forecasting is a popular task where the goal is to predict the short-term future trajectory of multiple agents in a traffic scene. Being able to predict the future positions and intents of each vehicle provides context for other modules in autonomous driving, such as path planning. Large, robust benchmarks such as the Waymo Motion Dataset [8, 9], Argoverse [10], and the NuScenes Dataset [11] have provided a standardized setting for advancements in the task, with leaderboards showing clear rankings for state-of-the-art models. Amongst the top performing architectures, most are based on Transformers for feature extraction [7, 12, 13, 14, 15]. Current SOTA models also model trajectory prediction probabilistically, as inspired by the use of GMMs in Multipath [16].
@@ V-B Performance on a Smaller Dataset Setting
We train the baseline model and all formulations on only 1% of the original Waymo dataset and benchmark their performance on 100% of the evaluation set in Table II. All experiments were trained over 50 epochs. In the small dataset setting, we observe that providing a kinematic prior in any form improves performance for minimum final displacement error (minFDE). Additionally, we observe better performance across all metrics for formulations 1, 2, and 1 with interpolation. Figure 4 shows how all kinematic formulations improve convergence speed over the baseline, with Formulation 4 (acceleration and steering) converging most quickly. Overall, Formulation 1 provides the greatest boost in mAP performance, with over $13\%$ gain over the vanilla baseline and $50\%$ over the deterministic baseline DKM, while Formulations 3 and 4 provide the greatest boost in distance metrics. In general, all formulations provide similar benefits in the small dataset setting, with the most general performance benefit coming from Formulation 3. While Formulation 4 also provides a comparable performance boost, the difference from Formulation 3 may be attributed to compounding error from second-order approximation. Interestingly, we note that modeling mean trajectories with kinematic priors is not enough to produce performance gain, as shown by the lower performance of DKM compared to vanilla MTR.
Compared to the results from Table I, the effects of kinematic priors in learning are more pronounced. Since kinematic priors naturally relate the position at one timestep to the position at the next, we believe that the performance boost can be attributed to this natural relationship. In backpropagation, optimization of one position further into the time horizon directly influences predicted positions at earlier timesteps via the kinematic model. Without the kinematic prior, the relation between timesteps may be implicitly related through neural network parameters. When the model lacks data to form a good model of how an agent moves through space, the kinematic model can help to compensate by modeling simple constraints.
@@ V-C Performance in the Presence of Noise
TABLE III: Percentage (%) degradation of performance in the presence of noise. We also compare the robustness of each method by measuring the impact on performance in the presence of noise. In the table, we compute the relative change in performance between perturbed evaluation and clean evaluation, relative to each method. We observe that the model trained with Formulation 4 (steering and acceleration representations) offers the greatest robustness over other formulations in the presence of noisy inputs.
@@ VI Discussion and Conclusion
When there is large-scale data to learn a good model of how vehicles move, we observe that the effects of kinematic priors are less pronounced. This is demonstrated by the less obvious improvements over the baseline in Table I compared to Table II; model complexity and dataset size will eventually out-scale the effects of the kinematic prior. With enough resources and high-quality data, trajectory forecasting models will learn to “reinvent the steering wheel”, or implicitly learn how vehicles move via the complexity of the neural network. For practical autonomous systems, portability and adaptability are important problems to consider aside into the future. We hope that kinematic priors can be considered in trajectory forecasting and simulation as deployment ventures past well-represented urban scenarios to suboptimal settings, whether it is with noisy sensors, smaller datasets to adapt from, or downscaled models for inference speed and portability.
One limitation of our work is that we primarily explore analysis in one-shot settings; future work focusing on kinematic priors for autoregressive approaches would be interesting in comparison to one-shot models with kinematic priors, especially since autoregressive approaches model predictions conditionally based on previous timesteps.